OpenStudy (anonymous):

Evaluate the integral ∫R2xydA, where R is the triangle 2x+y≤2, x≥0, y≥0

OpenStudy (anonymous):

hey mr. math

OpenStudy (anonymous):

u were there earlier with james right?

OpenStudy (mr.math):

Yes.

OpenStudy (anonymous):

okay...we came up with the integral...but it seems like my bounds are incorrect

OpenStudy (mr.math):

What were your bounds?

OpenStudy (anonymous):

we came up with x from 0 to 1...and y from 0 to -2x+2

OpenStudy (anonymous):

2xy dydx

OpenStudy (mr.math):

This is correct.

OpenStudy (mr.math):

This is correct.

OpenStudy (anonymous):

how come i can't run that through wolfram?

OpenStudy (mr.math):



OpenStudy (anonymous):

right we got all that...and looks the same...i try to run it through wolfram to get an exact answer and it doesn't work

OpenStudy (mr.math):

It should look like\[\int\limits_0^1 \int\limits_0^{2-2x} 2xydydx\]

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

lemme try that

OpenStudy (anonymous):

unreal...i kept getting -1/3....so i ran it through wolfram cause i thought there was an error....the answer is 1/3

OpenStudy (mr.math):

\[=\int\limits_0^1xy^2|_0^{2-2x}dx=\int\limits_0^1x(2-2x)^2=4\int\limits_0^1x^3-2x^2+x\] \[=(x^4-\frac{8}{3}x^3+2x^2)|_0^1=1-\frac{8}{3}+2=\frac{1}{3}.\]

OpenStudy (anonymous):

ya i got it now...thx soo much for helping

OpenStudy (jamesj):

now try the order of integration reversed